Jim's Farm's payables deferral period (PDP) is 50 days (on a 365-day basis), accounts payable are $100 million, and its balance sheet shows inventory of $154.3 million. What is the inventory turnover ratio?
Answer #1

payables deferral period =(Accounts payable/COGS)*365 days

50=100/COGS*365

(50/365)=100/COGS

Hence COGS=100/(50/365)_

=$730 million.

Inventory turnover ratio=COGS/Inventory

=730/154.3

=4.73 (Approx).